# Colimits of Heyting Algebras through Esakia Duality

### v1

@@ 6 Categories of Posets with P-morphisms
We now turn our attention to how the questions from the previous sections can be understood in a “discrete” setting – i.e., by considering posets, rather than Esakia spaces. One key motivation for this comes from the logical setting, where such discrete duals correspond to the “Kripke semantics” of modal and superintuitionistic logics. As we will see, a construction analogous to the $V_{G}$ functor can be provided, relating the category of Image-Finite Posets with p-morphisms, and the category of posets with monotone maps. To justify it, we begin by outlining the relationship between such a category and the category of finite posets with p-morphisms, and the general phenomenon of profiniteness. We then describe the full construction.
@@ 6.1 Pro- and Ind-Completions of categories of finite objects
To explain the questions at hand, we recall the picture, outlined in [19, Chapter VI], relating categories of finite algebras and several of their completions. To start with, we have $\mathbf{FinBA}$, the category of finite Boolean algebras, and its dual, $\mathbf{FinSet}$. For such categories, there are two natural “completions”:
@@
where $\mathbf{DL}$ is the category of all distributive lattices with lattice homomorphisms; $\mathbf{CCJDL}$ is the category of complete and completely join-prime generated distributive lattices with complete lattice homomorphisms; $\mathbf{Pos}$ is the category of all posets with monotone maps; and $\mathbf{Pries}$ is the category of Priestley spaces with Priestley morphisms.
If we are interested in Heyting algebras, it is not hard to see that the category of finite Heyting algebras with Heyting algebra homomorphisms is dual to the category $\mathbf{FinPos}_{p}$ of finite posets with p-morphisms. The Ind-completion of $\mathbf{FinHA}$ is of course still $\mathbf{HA}$, and what we have shown above is that the inclusion of $\mathbf{HA}$ into $\mathbf{DL}$ admits a left adjoint. So if we wish to use this abstract perspective to “guess” a discrete category, the natural place to look is the Ind-completion of $\mathbf{FinPos}_{p}$. Such objects are those which are directed colimits of finite posets via p-morphisms – but this means that the finite posets must embed into the top of the object, i.e., such objects must be image-finite posets, in other words
@@
where $\mathbf{ImFinPos}$ is the category of image-finite posets with p-morphisms. And indeed, we have that such a category is, by the main results from [4], dual to the category of profinite Heyting algebras, as desired. Indeed, a similar picture to this one has been discussed for finite modal algebras, and their duals, finite Kripke frames, in the recent preprint [13], where the above dualities are briefly sketched  [^3] Implicit in the discussion above is the Raney duality between complete and completely join-prime generated distributive lattices, and its restriction to Heyting algebras. Whilst this result is very well-known and often mentioned in the literature, and traces its origins to the work of Raney, De Jongh and Troelstra, a concise proof of it does not seem to be available in the literature; in the Appendix we provide such a proof, for completeness.
@@ 6.2 An Adjunction between the category of Image-Finite posets and Posets with Monotone Maps
We write $P_{G}^{g}(P_{\bullet}^{g}(X))$ for the image-finite part of the projective limit of the above sequence in the category of posets with monotone maps. In other words, we take
@@
Using this we can show that the poset $P_{G}^{g}(X)$ will be terminal in the desired way. The next Proposition contains the crux of the arguments:
@@
Using similar arguments, and similar ideas to Proposition 16, we can then show:
@@
A few facts are worthy of note here: following the main ideas of [13], the above describes an adjunction which splits the adjunction between profinite Heyting algebras and the category of sets, a fact which is worthy of note. Moreover, as noted in this work, the construction $P_{G}$ given above can be thought of as a generalization of the $n$-universal model – indeed, if one starts with the dual poset of the free distributive lattice on $n$-generators, $P_{G}$ will produce precisely this same model. This illustrates an interesting connection between these two well-known constructions of the free algebra – the Ghilardi/Urquhart step-by-step construction, and the Bellissima/Grigolia/Shehtman universal model – showing that they are, in some sense, dual to each other.
@@ 7 Conclusions and Further Research
In this note we have generalized Ghilardi’s construction of the free Heyting algebra generated by a finite distributive lattice to the general case, and have extracted some consequences from this – namely, the fact that the category of Heyting algebras is co-distributive, and direct proofs of amalgamation of Heyting algebras deriving from the corresponding properties for distributive lattices, together with general considerations. We have also looked at how this construction fares in different settings: when considering specific subvarieties of Heyting algebras (Boolean algebras, $\mathsf{KC}$-algebras, $\mathsf{LC}$-algebras), where it is shown that the construction can be reasonably adapted, and sometimes corresponds to natural modifications of the Ghilardi functor; and also we considered the relationship between the category of image-finite posets with p-morphisms and the category of posets with monotone maps, showing that reasonable modifications also yield an adjunction here.
Several questions are raised by the above considerations, both of a technical and of a conceptual nature. An immediate application which we intend to develop in a companion paper to this one is that the above adjunctions can be used to provide a coalgebraic semantics for some intuitionistic modal logics; this is done by showing that essentially the coalgebraic representation which works for distributive lattices can be “reflected” to Heyting algebras by using the $V_{G}$ and $P_{G}$ constructions.
Looking futher, in the case of finite distributive lattices, Ghilardi managed to show using his construction that free finitely generated Heyting algebras are bi-Heyting algebras; this relies on the fact that applying $V_{g}(-)$ to a finite poset yields a finite poset, which has a natural structure as a bi-Esakia space. It is not clear that even starting with a bi-Esakia space $X$, $V_{g}(X)$ will be bi-Esakia; we have been able to show that $V_{r}(X)$ will be co-Esakia, but all other facts remain open.
Following in the study of subvarieties of Heyting algebras, a systematic study connecting correspondence of axioms with Kripke semantics, and the appropriate modifications made to the $V_{G}$ functor, seems to be in order. This does not seem immediately straightforward, as our example with $\mathsf{KC}$ algebras illustrates, but it might be possible to obtain more general results.
@@
Finally, the present methods show that the category of image-finite posets includes into all posets through an inclusion admitting a right adjoint, but this explains nothing of the inclusion
$\mathbf{Pos}_{p}\to\mathbf{Pos}.$
The former category seems even less well-behaved than the category of image-finite posets, and seems to be connected with the category of Topological spaces with open and continuous maps. We leave it for further research to study how such adjunctions operate, and their consequences for the study of categories of logic, of algebras, their connections with point-free topology, and other related fields.
8 Acknowledgements
I would like to thank Silvio Ghilardi, Nick Bezhanishvili and Lingyuan Ye for several inspiring conversations and suggestions on parts of this work. I would also like to especially thank Nick Bezhanishvili for several suggestions of improvement on earlier versions of this work. I especially want to thank Mamuka Jibladze for many enthusiastic discussions on the main construction of this paper, and for his key idea in proving a lemma above.

### v3

@@ 6 Categories of Posets with P-morphisms
We now turn our attention to how the questions from the previous sections can be understood in a “discrete” setting – i.e., by considering posets, rather than Esakia spaces. This can be motivated from a categorical point of view, where it arises as a natural question concerning the existence of adjunctions in $\mathsf{Ind}$ and $\mathsf{Pro}$-completions of categories. Another motivation for this comes from the logical setting, where such discrete duals correspond to the “Kripke semantics” of modal and superintuitionistic logics. As we will see, a construction analogous to the $V_{G}$ functor can be provided, relating the category of Image-Finite Posets with p-morphisms, and the category of posets with monotone maps.
@@ 6.1 Pro- and Ind-Completions of categories of finite objects
To explain the questions at hand, we recall the picture, outlined in [johnstone1982stone, Chapter VI], relating categories of finite algebras and several of their completions; we refer the reader to [Massas2023] where these and other dualities are outlined. To start with, we have $\mathbf{FinBA}$, the category of finite Boolean algebras, and its dual, $\mathbf{FinSet}$. For such categories, there are two natural “completions”:
@@
where $\mathbf{DL}$ is the category of all distributive lattices with lattice homomorphisms; $\mathbf{CCJDL}$ is the category of complete and completely join-prime generated distributive lattices with complete lattice homomorphisms; $\mathbf{Pos}$ is the category of all posets with monotone maps.
If we are interested in Heyting algebras, it is not hard to see that the category of finite Heyting algebras with Heyting algebra homomorphisms is dual to the category $\mathbf{FinPos}_{p}$ of finite posets with p-morphisms. The Ind-completion of $\mathbf{FinHA}$ is not necessarily $\mathbf{HA}$, since not every Heyting algebra is a directed colimit of finite Heyting algebras; however, one may use this intuition to “guess” a discrete category which allows the generalization of the above results. Indeed, the natural place to look is the Ind-completion of $\mathbf{FinPos}_{p}$. Such objects are those which are directed colimits of finite posets via p-morphisms – but this means that the finite posets must embed into the top of the object, i.e., such objects must be image-finite posets, in other words
@@
where $\mathbf{ImFinPos}$ is the category of image-finite posets with p-morphisms. And indeed, we have that such a category is, by the main results from [Bezhanishvili2008], dual to the category of profinite Heyting algebras, as desired. Indeed, a similar picture to this one has been discussed for finite modal algebras, and their duals, finite Kripke frames, in the recent preprint [profinitenessmonadicityuniversalmodels], where the above dualities are briefly sketched.
@@ 6.2 An Adjunction between the category of Image-Finite posets and Posets with Monotone Maps
We write $P_{G}^{g}(X)$ for the image-finite part of the projective limit of the above sequence in the category of posets with monotone maps. In other words, we take
@@
Using this we can show that the poset $P_{G}(X)$ will be free in the desired way. The key fact that is missing concerns the fact that, after applying Lemma 57 infinitely often, one needs, as in Proposition 20, to show that the lifting is indeed a p-morphism. There we used compactness, which certainly does not hold for arbitrary posets.
@@
Using similar arguments as those of Theorem 22, we can then show:
@@
A few facts are worthy of note here: following the main ideas of [profinitenessmonadicityuniversalmodels], the above describes an adjunction which splits the adjunction between profinite Heyting algebras and the category of sets. Moreover, as noted in [profinitenessmonadicityuniversalmodels], the construction $P_{G}$ given above can be thought of as a generalization of the $n$-universal model – indeed, if one starts with the dual poset of the free distributive lattice on $n$-generators, $P_{G}$ will produce precisely this same model. This illustrates an interesting connection between these two well-known constructions of the free algebra – the Ghilardi/Urquhart step-by-step construction, and the Bellissima/Grigolia/Shehtman universal model – showing that they are, in some sense, dual to each other.
It is also notable that the above methods cannot be expected to be extended to yield an adjunction between $\mathbf{Pos}_{p}$ and $\mathbf{Pos}$ as above. Indeed, as noted in [openmapsdoesnothaveproducts], $\mathbf{Pos}_{p}$ does not have binary products. If $I:\mathbf{Pos}_{p}\to\mathbf{Pos}$ had a right adjoint, then all limits in $\mathbf{Pos}$ would have to be preserved, which is absurd.
@@ 7 Conclusions and Further Research
In this paper we have explicitly described the adjunction between Heyting algebras and Distributive lattices, and have extracted some consequences for the theory of Heyting algebras from this – namely, facts about the structure of free Heyting algebras, properties of the category of Heyting algebras, and direct proofs of amalgamation of Heyting algebras deriving from the corresponding properties for distributive lattices. We have also looked at how this construction fares in different settings: when considering specific subvarieties of Heyting algebras (Boolean algebras, $\mathsf{KC}$-algebras, $\mathsf{LC}$-algebras), where it is shown that the construction can be reasonably adapted, and sometimes corresponds to natural modifications of the Ghilardi functor; and also we considered the relationship between the category of image-finite posets with p-morphisms and the category of posets with monotone maps, showing that reasonable modifications also yield an adjunction here.
As noted throughout the text, the material presented here does not exhaust the range of applications of these constructions and the ideas wherein: in [openmapsandinterpolantcomplexity] we expect to present proofs of uniform interpolation, as well as new results concerning interpolant complexity, and the structure of free Heyting algebras, exploiting the description of coproducts given above. Moreover, this research leaves open several questions, both of a technical and of a conceptual nature.
One fact which we have not worked with concerns the structure of the Priestley spaces obtained by applying $V_{g}(-)$; when $X$ i a finite poset, this yields a finite poset, which has a natural structure as a (bi-)Esakia space. It is not clear that even starting with a bi-Esakia space $X$, $V_{g}(X)$ will be bi-Esakia; we have been able to show that $V_{r}(X)$ will be co-Esakia, but all other facts remain open.
Following in the study of subvarieties of Heyting algebras, a systematic study connecting correspondence of axioms with Kripke semantics, and the appropriate modifications made to the $V_{G}$ functor, seems to be in order. This does not seem immediately straightforward, as our example with $\mathsf{KC}$ algebras illustrates, but it might be possible to obtain more general results, at least in the case of finitely axiomatisable such logics.
@@
Finally, the present methods open up the possibility of a more combinatorially rich analysis of Heyting algebras. A first question of this type, raised above, concerns whether the $V_{G}$-construction admits any fixed points, i.e., any Esakia spaces which are isomorphic to their free extension. Several related questions can be asked, which might be of interest both to the intrinsic study of Heyting algebras and in applications to intuitionistic forcing and intuitionistic mathematics, given they comprise the combinatorial fine-structure of duals of Heyting algebras. We leave this as an exciting line of research to be pursued.
[…]
[ heading=bibintoc, title=Bibliography ]